OpenStudy (anonymous):

how to integrate (dy/dx)=[3(x+y)] /x ?

OpenStudy (nincompoop):

distributive property then separate them as terms with denominator x

OpenStudy (anonymous):

=3+(3y/x)???

OpenStudy (nincompoop):

distributive property

OpenStudy (anonymous):

yes, so distribute the 3 and then split up the numerator~?

OpenStudy (nincompoop):

have a rest and do it again when you have a refreshed brain :P

OpenStudy (anonymous):

? wait wat?

OpenStudy (nincompoop):

distributive property is multiplication yes? 3(x+y) = ?

OpenStudy (anonymous):

3x+3y, but then there is a x on the bottom, so then the x's cancel in the first term, leaving just 3, and the second term is left with 3y/x.

OpenStudy (nincompoop):

1/x * (3x + 3y)

OpenStudy (nincompoop):

okay then you have integral (3 + 3xy) dx

ganeshie8 (ganeshie8):

looks its not separable, familiar with computing IF and multiplying both sides... ?

OpenStudy (anonymous):

its a homogenous equation try using y=vx so dy/dx=v+xdv/dx

OpenStudy (nincompoop):

integral (3 + 3x/y) dx

OpenStudy (anonymous):

???????? confusion!!!! its not seperable?

OpenStudy (anonymous):